OpenStudy (vuriffy):

The variables x and y satisfy the equation 5^(y+1) = 2^(3x) By taking logarithms, show that the graph of y against x is a straight line. ii) Find the exact value of the gradient of the line and state the coordinates of the point at whch the line cuts the y-axis.

OpenStudy (vuriffy):

@Hayhayz Another one she didn't teach us :o

OpenStudy (3mar):

If I may help?

OpenStudy (vuriffy):

If you please.

OpenStudy (3mar):

With my pleasure!

OpenStudy (3mar):

Can you apply log to both sides with any base you want?

OpenStudy (vuriffy):

I have no idea..

OpenStudy (3mar):

\[\Large 5^{(y+1)} = 2^{(3x)}\] Is that your equation?

OpenStudy (vuriffy):

Yes.

OpenStudy (3mar):

Nice!

Hayhayz (hayhayz):

you can use logs to help you here

OpenStudy (vuriffy):

Do I like ln both sides to move the exponents to the front?

OpenStudy (3mar):

Take log (standard log with base of 10) for both side, you get: \[\Large \log[5^{(y+1)}] = \log[2^{(3x)}]\]

Hayhayz (hayhayz):

yup u got it

Hayhayz (hayhayz):

remember log(A^Y) = y log(A)

OpenStudy (3mar):

And one of the properties of logarithm's is \[\huge \log(x^a)=alogx\]

OpenStudy (vuriffy):

I got: (y + 1)log (5) = (3x)log (2)

OpenStudy (3mar):

Great!

Hayhayz (hayhayz):

if you proved log(A*B) = log(A)+LOG(B) then you can easily see that log(A^N) = log(A*A*A*A*A..) = LOG(A) + LOG(A) + LOG(A)+... = n log (A)

OpenStudy (3mar):

Thank you for the medal! But let's finish it firstly

OpenStudy (vuriffy):

or y = ?

OpenStudy (3mar):

\[\Large (y+1)*\log[5] = (3x)\log[2]\] then \[\Large (y+1)= (3x)\frac{ \log[2] }{ \log[5] }\] then \[\Large (y+1)= const*(3x)\] then \[\Large y+1= const*x\] then \[\Huge y= const*x-1\] which is an equation of straight line

OpenStudy (3mar):

I hope I helped you.

OpenStudy (vuriffy):

Wonderful, I got that, but didn't put the "constant" part.

OpenStudy (3mar):

Thank you for your appreciation! You can calculate it easily I think!

OpenStudy (vuriffy):

I did 3x - log(2)/log(5) - 1 3x - 0.5693 and so on.

Hayhayz (hayhayz):

algebra looks wrong

OpenStudy (3mar):

log(2)/log(5)=?

OpenStudy (vuriffy):

0.43067?

OpenStudy (3mar):

0.43067*3=?

OpenStudy (vuriffy):

AH

OpenStudy (3mar):

I think you got it!

OpenStudy (vuriffy):

y = 3x - 0.292?

OpenStudy (3mar):

y=(0.43067*3)x-1

OpenStudy (vuriffy):

Oh, I get it now, that's what confused me.

OpenStudy (vuriffy):

How would you find the gradient, isn't it the m value?

OpenStudy (vuriffy):

1.292?

OpenStudy (3mar):

Yes it is the slope/the gradient = 1.292 Brilliant!

OpenStudy (vuriffy):

So for the y-intercept, all I would have to do is plug in 0 for x and get y = -1? The cords would be (0, -1) ?

OpenStudy (3mar):

Excellent! I think you got it!

OpenStudy (vuriffy):

Wonderful, thank you so much.

OpenStudy (3mar):

You are welcome! Any Help... Any Time... Thank you for learning!

OpenStudy (3mar):

Do you have any more questions?

OpenStudy (vuriffy):

Actually, there's one in my notes I never finished. Want me to tag you?

OpenStudy (3mar):

If you need a help and if I am not bothering you!

OpenStudy (vuriffy):

Do you mind if I just type in out right now?

OpenStudy (vuriffy):

It is given that ln(y+1) - lny = 1 +3lnx. Express y in terms of x, in a form not involving logarithms.

OpenStudy (3mar):

Can we start in new post, please?

OpenStudy (vuriffy):

Yes, of course.

OpenStudy (3mar):

Welcome!